Find the specific gravity and the density of an object with a dimension of 25cm x 60cm x 80cm with a mass of 9000g.


1
Expert's answer
2021-02-12T06:10:51-0500

Density:

ρ=mVρ=9000256080=0.075gcm3\rho=\frac{m}{V}\\\rho=\frac{9000}{25\cdot60\cdot80}=0.075\frac{g}{cm^3}

The specific gravity:


SG=ρρw=0.0751=0.075SG=\frac{\rho}{\rho_w}=\frac{0.075}{1}=0.075